### v1

# arXiv:2603.00774v1[cs.HC] 28 Feb 2026

Abstract.
While large language models (LLMs) excel at open-ended dialogue, effective psychotherapy requires structured progression and adherence to clinical protocols, making the design of psychotherapist chatbots challenging. We investigate how different LLM-based designs shape perceived therapeutic dialogue in a chatbot grounded in the Self-Attachment Technique (SAT), a novel self-administered psychotherapy rooted in attachment theory. We compare three architectural variants: (1) a multi-agent system utilizing finite state machine aligned with therapeutic stages and a shared long-term memory, (2) a single-agent using identical knowledge-base and the same prompts, and (3) an unguided LLM. In an eight-day randomized controlled trial (RCT) with $N=66$ Farsi-speaking participants, balanced across the three chatbots, the multi-agent system is perceived as significantly more natural and human-like than the other variants and achieves higher ratings across most other metrics. These findings demonstrate that for therapeutic AI, architectural orchestration is as critical as prompt engineering in fostering natural, engaging dialogue.
Therapeutic chatbots, LLMs, conversational agents, digital mental health, multi-agent systems
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfacesccs: Applied computing Health care information systems
Figure 1. Overview of the user study comprising three phases: (1) recruitment and blinded RCT group assignment; (2) an eight-day study period during which participants interacted with one of three therapeutic chatbot versions, multi-agent FSM-based, single-agent with therapy knowledge, or unguided single-agent; and (3) anonymized post-study survey completion data analysis.
1. Introduction
Conversational agents are increasingly explored as scalable mental health support tools, offering on-demand access, anonymity, and reduced cost compared to traditional care (Jabir et al., 2023; Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Advances in large language models (LLMs), such as ChatGPT, now enable open-ended, context-sensitive conversations, expanding the expressive capacity of mental health chatbots (Lee et al., 2020; Moell, 2024). At the same time, structured therapeutic frameworks can be embedded into conversational systems. The Self-Attachment Technique (SAT) (Edalat, 2015) is a self-administered, attachment-based approach that supports emotion regulation and self-compassion by cultivating an imaginative affectional bond between the adult self and the childhood self of an individual (Alazraki et al., 2021; Elahimanesh et al., 2023). Prior SAT chatbots using rule-based or classification-driven dialogue demonstrated promising empathy and engagement, but were limited in flexibility and scalability (Alazraki et al., 2021; Elahimanesh et al., 2023).
Despite these advances, most mental health chatbots emphasize feasibility and acceptability while relying on shallow interaction patterns and short-term evaluations, leaving open questions about how system architecture shapes perceived therapeutic quality (Kang and Hong, 2025; Tong et al., 2025; He et al., 2023a). Research highlights the importance of personalization, empathic responses, and sustained engagement (He et al., 2023a; Jabir et al., 2023), yet existing work rarely compares architectural paradigms under controlled conditions (Zhang et al., 2025). Recent proposals such as multi-agent orchestration and finite state machines offer ways to structure complex LLM behavior (Zhang et al., 2025), but their impact on therapeutic naturalness, authenticity, and engagement, particularly for structured interventions like SAT, remains underexplored.
To address these gaps, we conducted a between-subjects randomized controlled trial (RCT) examining how LLM-based architectural design influences perceived therapeutic quality in a SAT-informed chatbot. Using GPT-4o (OpenAI et al., 2024) as the base model across all conditions, we compared three architectures. The Intervention employed a multi-agent finite state machine aligned with SAT stages, incorporating explicit state transitions, a SAT knowledge base, and shared long-term memory among the agents. The Active Control used a single agent with equivalent prompts and SAT access. The Placebo Control used GPT-4o without SAT knowledge or structure, resembling a generic conversational companion. System architecture and prompts were authored in English, while deployment targeted a Farsi-speaking user population.
We deployed the system for eight days with $N=66$ participants (Intervention: 22, Control: 23, Placebo: 21), who completed demographic measures, interacted with their assigned chatbot, and completed post-study surveys. The intervention was perceived as significantly more natural and human-like than both control conditions (F=7.017, p=0.0018, $\eta^{2}$=0.187), with mostly higher but non-significant scores on other therapeutic metrics, such as trust, memory coherence, user satisfaction, fewer off-topic responses, and ease of use. Chat logs showed the intervention produced more but shorter messages (459 messages, $\sim$230 characters) than the active control (336 messages, $\sim$409 characters) and placebo control (206 messages, $\sim$635 characters), indicating a more interactive dialogue style. Findings demonstrate that architectural choices, beyond prompt design alone, meaningfully shape user experience in LLM-based therapeutic chatbots.
2. Related Work
The Self-Attachment Technique (Edalat, 2015, 2016), grounded in Bowlby’s attachment theory (Bowlby, 2010, 2008), helps users regulate emotional distress by alternating between care-seeking and care-giving roles. Early SAT chatbots operationalize this with rule-based dialogue and emotion classification to provide empathetic responses (Alazraki et al., 2021; Law et al., 2022; Edalat et al., 2025), enforcing structure but remaining limited in generative flexibility, contextual adaptation, long-term personalization, and multilingual support. Similarly, other therapeutic chatbots, including CBT systems (Jabir et al., 2023; He et al., 2023a), Evebot (Yin et al., 2019), GPT-2–based therapy bots (Wang et al., 2021), and student distress-detection frameworks (Patel et al., 2019), combine emotion recognition with neural generation but generally lack explicit therapeutic state modeling, memory, and robustness. The emergence of large language models has enabled richer human–AI collaboration and motivated hybrid neuro-symbolic architectures. Human-in-the-loop systems support cognitive restructuring (Sharma et al., 2024), yet CAPE-II evaluations report failures in crisis handling and boundary maintenance (Sobowale et al., 2025), and excessive trust can amplify harm (Song et al., 2025; Choi et al., 2025). FSM-based systems like MindfulDiary (Kim et al., 2024) and ChaCha (Seo et al., 2024) enforce structured progression while allowing free expression, multi-agent frameworks improve robustness and empathy (Zhang et al., 2025; Hu et al., 2025), EmoAgent monitors user state to reduce deterioration (Qiu et al., 2025), and LLMs serve as scalable front-ends for structured intake and assessment (Rosteck et al., 2025). Compared to early SAT bots (Alazraki et al., 2021; Elahimanesh et al., 2023), these architectures provide stronger guarantees of consistency, safety, and flexibility.
Emotion modeling and personalization are central to therapeutic effectiveness, optimizing empathy, fluency, and novelty (Alazraki, 2021), adapting interventions via real-time mood detection (EMMA; (Ghandeharioun et al., 2019)), and strengthening therapeutic bonds (Kocielnik et al., 2019). RCTs validate their impact: Woebot (Fitzpatrick et al., 2017a) reduced symptoms compared to self-help, while modality and interaction style influence loneliness and emotional dependence (Fang et al., 2025); other studies highlight emotional resilience (Ly et al., 2017), humor (Sun et al., 2024), and just-in-time interventions (Ghandeharioun et al., 2019). Farsi-language systems remain scarce: early Persian SAT chatbots report positive engagement (Elahimanesh et al., 2023), while HamRaz and HamRazEval support culturally grounded therapy (Abbasi et al., 2025) and PersianLLaMA provides a foundation for generative modeling (Abbasi et al., 2023), though personalization and fluency remain limited. Our work synthesizes these strands by embedding a context-aware LLM within a multi-agent, FSM-driven SAT architecture, advancing prior SAT and Persian systems with structured progression, safety-aware generation, long-term engagement, and one of the first RCT-based evaluations of a Farsi SAT chatbot.
3. Method
System Design and Architecture. To investigate how conversational structure and adaptivity influence therapeutic engagement, we developed a three-tier conversational AI system that delivers the Self-Attachment Technique within a Randomized Controlled Trial framework. The study included three experimental conditions: (1) a finite state machine (FSM) driven multi-agent system with long-term memory intervention (Alpha), (2) a single-agent system with a knowledge-base as an active control (Beta), and (3) a minimal single-agent placebo (Gamma). All versions run on a Django REST backend in Python and employ OpenAI’s GPT-4o (OpenAI et al., 2024) as the conversational engine. User interactions occurred in Farsi, supported by a specialized response retrieval system for intent classification, while prompts and conversational structures were authored in English. To maximize ecological validity, participants accessed the intervention via a React.js web interface, identical across all three conditions.
Experimental Conditions and Randomization. Upon registration, participants were randomly assigned to one of three groups and blinded to their condition to minimize expectancy effects. All systems employed the same LLM and shared interface, response style, and interaction modality. (1) Alpha implements a 12-state therapeutic protocol based on the SAT stages (Elahimanesh et al., 2023), while supporting open-ended conversation throughout the flow. At each stage, dedicated agents guide state transitions using LLM-driven decision logic, while maintaining open-ended dialogue to engage users and smoothly navigate the conversation to the next stage. To ensure consistency among the agents, the system utilizes a shared long-term conversational memory, generated via prompt-guided LLM summarization to preserve emotional states, personal context, and salient experiences. Additionally, it incorporates an agent using an adaptive Retrieval-Augmented Generation (RAG) mechanism to deliver personalized therapeutic exercises; (2) Beta delivers the identical SAT content and daily exercises but without FSM enforcement, relying solely on a single comprehensive system prompt to implicitly guide the conversation through similar phases within the LLM’s context window. To isolate the effect of architecture, this version uses the exact same prompts as Alpha, collapsed into a single sequential prompt; (3) Gamma provides a minimal conversational single-agent, controlling for the “digital companion” effect. It uses a stripped-down prompt with no SAT knowledge, no exercise recommendation engine, and no structured therapeutic goals, representing a near–pure instantiation of the LLM.
Conversational Structure: The Finite State Machine. A key contribution of the Alpha system is its explicit modeling of the therapeutic process. The FSM divides the interaction into four logical Superstates or phases: Initiation (Greeting/Emotion Assessment), Exploration (Event Processing), Intervention (Exercise Delivery), and Conclusion. Unlike rigid rule-based chatbots that rely on keyword matching, the Alpha system employs two advanced transition mechanisms to maintain naturalness: (1) LLM-as-Judge Sufficiency Detection (Zheng et al., 2023a; Szymanski et al., 2025), where a dedicated prompt evaluates whether accumulated user responses satisfy emotional and informational criteria for the current state, ensuring meaningful expression of the feelings and events before state change; (2) Intent-Based Routing, which classifies over 40 affirmative/negative Farsi linguistic patterns to route the user to the correct next state.
Knowledge Injection and Adaptive RAG. A comprehensive SAT knowledge base, containing theoretical foundations and 27 exercises, was embedded into the system prompts for both Alpha and Beta. Delivery, however, differed: Beta used a static schedule, selecting exercises from a fixed pool based on the agent, while Alpha employed a hierarchical Adaptive RAG mechanism. This process consists of two phases: (1) Constraint-Based Filtering: The system first retrieves a valid candidate set by filtering the exercise database according to the user’s protocol timeline (Intervention Day 1-8) and therapeutic stage (Beginning, Intermediate, or Advanced). (2) Semantic Selection: A ’Selector’ LLM then acts as a re-ranker, analyzing the user’s current emotional state and summarized memory to choose the single most contextually appropriate exercise from the candidate set. It then generates a personalized version of that exercise tailored to the user’s context, minimizing repetition and maximizing personalization.
Longitudinal Memory and Progress Tracking. Alpha maintains structured long-term memory, creating concise summaries every three messages via prompt-guided LLM abstraction to preserve emotional patterns, preferences, and biographical context. A calendar-based model decouples timeline from interaction frequency, ensuring participants resume the correct stage (capped at Day 8) and follow the protocol consistently from beginner to advanced levels.
4. Experiment and Results
Participant Recruitment, Demographics, Study Details. We conducted an eight-day study, following the eight phases of the SAT protocol (Edalat, 2015), with $N=66$ participants ($43$ male, $23$ female; mean age $27.2$, STD=$10.1$, range 1717-$73$), comparing three system variants: Alpha ($n=22$), Beta ($n=23$) and Gamma ($n=21$). Ethical approval for the user study was obtained from the institution of one of the authors. Participants were recruited by word of mouth and also from Farsi-speaking communities via social media (Twitter, Telegram), participated voluntarily, provided informed consent, and were informed that the system was not a replacement for professional mental health care; logs and survey responses were anonymized. Baseline measures of anhedonia, sadness, anxiety, and uncontrollable worry showed negligible effect sizes ($\eta^{2}\leq 0.053$), indicating balanced pre-existing distress. Most participants (47/66, 71.21%) reported high or very high familiarity with LLMs, while 49 (74.24%) had no prior familiarity with SAT, suggesting observed effects reflect system design rather than prior knowledge. The post-study survey used 5-point Likert items, assessing metrics adapted from prior mental health chatbot evaluations (Reguera-Gómez et al., 2025; Gao et al., 2025; Fitzpatrick et al., 2017b; He et al., 2023b). Statistical significance was assessed using one-way ANOVA (St et al., 1989) with $p$-values from 5,000-iteration permutation tests, a nonparametric approach that does not assume normality.
The Primacy of Conversational Naturalness. The strongest finding was a significant architectural effect on perceived naturalness and human-like tone. Alpha was rated higher ($\mu=3.955$, $SD=0.950$) than Beta ($\mu=3.043$, $SD=0.825$) and Gamma ($\mu=3.211$, $SD=0.787$) ($F=7.017$, $p_{perm}=0.0018$, $\eta^{2}=0.187$), reflecting a nearly one-point gain on a 5-point scale (0.912 over Beta) and a shift from “somewhat” to “quite” natural. The multi-agent FSM architecture explained $\sim$19% of rating variance, a moderate-to-large effect. Pairwise tests showed Alpha outperformed both Beta and Gamma, which were similar, indicating benefits of explicit multi-agent state management beyond content alone. All conditions showed comparable baseline usability (ease of use: $F=0.204$, $p=0.826$, $\eta^{2}=0.007$; feature integration: $F=0.859$, $p=0.446$, $\eta^{2}=0.027$), isolating improvements to relational quality rather than functional UX.
Table 1. Survey results compared three architectures on seven interaction metrics (1–5 Likert). Alpha almost consistently outperformed Beta and Gamma, with higher mean values. For off-topic replies, lower scores are better. Naturalness was statistically significant. F-statistics show variance between versus within groups, and $\eta^{2}$ indicates the proportion of variance explained by architecture.
Metric Alpha Beta Gamma $F$-statistic $\eta^{2}$
Naturalness* 3.955 3.043 3.211 7.017 0.187
Trust 3.136 2.739 2.737 0.737 0.024
Empathy 3.727 3.435 3.895 1.357 0.043
Memory 3.682 3.348 3.263 1.061 0.034
Satisfaction 3.636 3.348 3.474 0.520 0.017
Off-topic Replies 1.727 2.304 2.000 1.481 0.046
Ease of Use 4.000 3.826 3.789 0.204 0.007
Evaluated Interaction Metrics. As shown in Table 1, Alpha trended higher in the majority of other experiential metrics as well, including memory for past conversations ($\mu=3.682$ vs. 3.348/3.263), perceived empathy ($\mu=3.727$ vs. 3.435/3.895), and trust, while these differences were not statistically significant. Alpha also reported fewer off-topic replies ($\mu=1.727$ vs. 2.304/2.000), indicating better conversational focus, though not significant ($F=1.481$, $p=0.225$, $\eta^{2}=0.046$). Overall satisfaction was moderately positive across groups ($\mu\approx 3.4$–3.6, $F=0.520$, $p=0.606$, $\eta^{2}=0.017$), showing general acceptability regardless of architecture. These patterns suggest that architectural design mainly drives perceived naturalness and conversational dynamics, while also enhancing relational qualities like trust and empathy; detecting significance for these secondary effects may require larger samples or distinct interventions.
Interaction Dynamics: Why Architecture Matters. To understand why the multi-agent system achieved higher perceived naturalness, we analyzed chat logs. One key factor appears to be turn-taking, pacing, and message length. Normalized by group size, Alpha participants generated the most agent messages per person ($\approx 20.9$) versus Beta ($\approx 14.6$) and Gamma ($\approx 9.8$), a 43% and 113% increase. Alpha’s agent messages were shorter (229.8 characters) than Beta (408.7) and Gamma (635.1), shifting from long, lecture-style responses to rapid, dialogic exchanges that closely emulate human therapeutic conversation. User behavior mirrored this pattern: Alpha participants sent shorter messages on average (29.0 characters) than Beta (38.9) and Gamma (42.8), and the agent-to-user message length ratio was lowest in Alpha (7.9:1) versus Beta (10.5:1) and Gamma (13.4:1), indicating a more balanced, interactive dynamic.
In addition, importantly the FSM architecture in Alpha, explicitly structured around therapeutic stages, naturally reflects the reasoning of a human therapist, guiding the conversation through problem identification, exploration, and resolution. Multi-agent coordination keeps each agent focused on its stage, while long-term memory preserves context within and across sessions, enhancing continuity and relevance. A structured therapeutic knowledge base ensures contextually appropriate guidance. Beta’s prompting and knowledge contributed to its intermediate performance, and Gamma’s minimal design performed lowest. Together, these elements suggest that Alpha’s design, by mimicking human conversational dynamics, multi-agent specialization, memory retention, and grounded expertise, directly drives higher perceived naturalness and overall interaction quality.
5. Discussion
Our findings show that architectural choices in LLM-based therapeutic chatbots strongly shape user experience, beyond prompt engineering and toward fundamental questions of system structure, due to the staged, process-driven nature of therapy. Multi-agent architectures with explicit state management produced more natural interactions than single-agent or unguided LLMs, indicating that effective chatbots require careful orchestration alongside content delivery.
Conversational Naturalness. The intervention’s advantage in perceived naturalness reflects more than stylistic preference. Natural dialogue aligns with person-centered care, where interaction quality predicts outcomes (Miller and Rollnick, 2012). Chat logs show that the multi-agent system produced shorter, more interactive exchanges, shifting from monologic responses to genuine dialogue and challenging assumptions that chatbots should maximize information per turn. Instead, effective conversational AI should mirror human dialogue rhythms through frequent turn-taking and collaborative meaning-making. By distributing therapeutic stages across specialized agents, the architecture likely enabled this dialogic pattern, with long-term memory further contributing by preserving users’ feelings and experiences across sessions. Notably, gains in naturalness did not extend to empathy, trust, or personalization, suggesting that these are related yet distinct dimensions of the therapeutic alliance. Architectural improvements that enhance naturalness alone may therefore be insufficient to foster emotional attunement, which may depend more strongly on prompt design or response content.
Structural Scaffolding and Design Implications. Across most interaction metrics, including naturalness, trust, memory coherence, user satisfaction, off-topic responses, and ease of use, performance consistently followed Alpha >> Beta >> Gamma. This pattern highlights the role of therapeutic structure and architectural scaffolding in shaping perceived interaction quality, even without user awareness of system design. Stronger structural guidance supports coherent dialogue, relevant questioning, and alignment with therapeutic goals, without limiting conversational flexibility.
Limitations and Future Directions. Our study highlights several limitations and directions for future work: (1) the eight-day deployment captured only early user impressions, leaving long-term engagement, outcomes, and the development of user-agent relationships unresolved, including whether perceived naturalness translates into sustained use or longer-term benefits; (2) the use of single-item measures for empathy, trust, and personalization limits construct validity, and future studies should employ validated multi-item instruments (e.g., the Working Alliance Inventory (Hatcher and Gillaspy, 2006)) and examine how architectural effects vary across user characteristics; (3) the focus on Persian-speaking, highly educated participants (93.9% holding at least a Bachelor’s degree) and an English-authored system architecture constrains cross-linguistic and broader population generalizability; (4) therapeutic outcomes and skill acquisition were not directly assessed, underscoring the need to link subjective interaction quality to objective behavioral and clinical measures; and (5) the collected chatbot interaction data represent a valuable therapeutic dataset that, when combined with other sources, could support LLM alignment using methods such as PPO or DPO (Zheng et al., 2023b; Rafailov et al., 2024).
6. Conclusion
Architectural design strongly shaped user experience in our therapeutic LLM chatbot. Multi-agent, memory-augmented systems with explicit state management produced more natural, human-like interactions than single-agent or unguided alternatives, with higher (though non-significant) relational ratings. By holding content constant, we isolated structural effects often conflated with content. These results suggest that advancing therapeutic AI requires not only capable LLMs and refined prompts, but also careful architectural orchestration to support structured therapeutic progression.
Appendices
Appendix A Finite State Machine Specification
The Alpha (Intervention) system is operated using a programmatic Finite State Machine, comprising 12 distinct states, based on the flowchart of the SAT [12] algorithm. This architecture enforces the therapeutic structure of the Self-Attachment Technique, while maintaining conversational flexibility through LLM decision nodes. The FSM is divided into four logical phases: Initiation, Exploration, Intervention, and Conclusion.
A.1. Transition Mechanisms
State transitions are governed by two complementary mechanisms designed to balance protocol adherence with flexible, naturalistic interaction.
  1. (1)
    LLM-as-Judge Sufficiency Detection. For conversational states that require qualitative and experiential information (such as Emotion Assessment, Event Exploration), we adopt an LLM-as-judge paradigm. In this setup, a dedicated judging prompt instructs the LLM to evaluate whether the accumulated conversation history sufficiently satisfies the informational and emotional criteria defined for the current state. Transitions occur only when the LLM-as-judge determines that the user has meaningfully articulated their emotional state or relevant experiences, enabling adaptive progression without reliance on fixed turn counts.
  2. (2)
    Intent Classification via Fine-Tuned BERT. For decision points requiring binary or categorical user input (such as accepting or declining an exercise), we employ an intent classification module based on BERT [8]. The model was fine-tuned on Farsi-language interaction data to recognize user intent, including over 40 affirmative and 13 negative linguistic patterns (such as “bale”, “are”, “begoo”). This approach enables robust intent detection while avoiding rigid keyword matching, thereby supporting fluent and user-driven conversational flow.
A.2. Detailed State Descriptions
A.2.1. Phase I: Initiation and Assessment
  • State 1: GREETING_FORMALITY_NAME Objective: Establish rapport and configure user preferences. Logic: The agent greets the user and optionally requests their name and formality preference (formal vs. informal). The transition occurs once the Judge confirms these parameters are set or explicitly declined by the user.
  • State 2: EMOTION Objective: Identify the user’s current emotional state. Logic: The agent employs active listening to validate feelings. A strictly enforced constraint requires a minimum of two user messages in this state to prevent premature transitions, ensuring the user feels heard before moving forward.
  • State 3: EMOTION_DECIDER (Internal Node) Objective: Route the conversation based on sentiment analysis. Logic: An LLM classifier analyzes the dialogue history to categorize the user’s state as Positive or Negative.
    • Negative: Transitions to SUPER_STATE_EVENT for processing triggers.
    • Positive: Transitions directly to ASK_EXERCISE, skipping the event exploration phase.
A.2.2. Phase II: Exploration (Trigger Processing)
  • State 4: SUPER_STATE_EVENT Objective: Explore the antecedents of negative emotions. Logic: Accessible only via the Negative path. The agent probes for specific events or triggers associated with the user’s distress. Similar to the Emotion state, a minimum two-message constraint is enforced to ensure adequate exploration depth.
  • State 5: OPEN_ENDED_CONVERSATION Objective: Provide a safe space for deep elaboration. Logic: If the user indicates a need to vent or elaborate further during State 4, the system transitions here. This state allows for free-form dialogue (minimum 4 messages) before gently steering the user back to the intervention track.
A.2.3. Phase III: Therapeutic Intervention
  • State 6: ASK_EXERCISE Objective: Obtain informed consent for the intervention. Logic: The agent proposes a therapeutic exercise. The Response Retriever classifies the user’s reply; affirmative responses lead to suggestion, while negative responses route to the closing phase.
  • State 7: EXERCISE_SUGGESTION Objective: Deliver a context-aware SAT exercise. Logic: This state utilizes the RAG system to select an exercise. The selection algorithm filters by the user’s specific intervention day (such as Day 1: Exercises 1-3) and therapeutic stage (Beginning/Intermediate/Advanced), then uses the LLM to choose the specific activity that best aligns with the user’s summarized emotional context.
  • State 8: EXERCISE_EXPLANATION Objective: Guide the user through the selected practice. Logic: The system provides step-by-step instructions. The user can request clarification, proceed to feedback, or request a different exercise (looping back to State 10).
  • State 9: FEEDBACK Objective: Debrief and consolidate learning. Logic: The agent asks the user to reflect on their experience immediately post-exercise.
  • State 10: LIKE_ANOTHER_EXERCISE Objective: Iterative intervention loop. Logic: The system offers additional exercises. An affirmative response loops the FSM back to EXERCISE_SUGGESTION, allowing for multi-exercise sessions; a negative response transitions to termination.
A.2.4. Phase IV: Conclusion
  • State 11: THANKS Objective: Session closure and positive reinforcement. Logic: The agent provides concluding remarks and encouragement.
  • State 12: END Objective: Terminal state. Logic: The session is formally closed, and the final memory summary is committed to the long-term storage database.
Appendix B Screenshot of the Web-based Interface
Figure 2. Screenshot of the web-based user interface of the chatbot. After logging in, users are directed to the home screen where they can start interacting with the chatbot. (A) shows the list of user messages and corresponding chatbot responses. (B) is the input area for composing and sending messages to the chatbot. (C) contains the logout button on the left and the button to restart the conversation on the right.
Appendix C Extended Quantitative Analysis
This section provides a comprehensive breakdown of the statistical results from the post-study survey ($N=66$). While the main text focuses on the primary outcome of naturalness, we present here the full statistical profile for all measured dimensions to ensure transparency and support future replication.
C.1. Statistical Methodology
Given the sample size ($N=66$) and the potential for non-normal distribution in Likert-scale responses, we employed a non-parametric permutation testing approach for all One-Way ANOVA comparisons.
  • Resampling Strategy: For each metric, group labels were randomly shuffled 5,000 times to generate a null distribution of F-statistics.
  • Significance Thresholds: We report the permutation $p$-value ($p_{perm}$). Significance is defined at $\alpha<0.05$.
  • Effect Size: We report $\eta^{2}$ (Eta-squared) to quantify the proportion of variance explained by the architectural condition.
C.2. Comprehensive Survey Results
Table 2 details the descriptive statistics (Mean and Standard Deviation) and inferential statistics (F-value, $p$-value, $\eta^{2}$) for all interaction quality items. The results confirm one statistically significant effect: Naturalness ($p=0.0018$). All other differences did not reach statistical significance ($p>0.05$).
Table 2. Complete Statistical Analysis of User Experience Survey Items ($N=66$). Significance is determined via 5,000-iteration permutation testing. Bold indicates the maximum mean per row; significance is reported separately ($\alpha<0.05$).
Survey Item Alpha (Int.) Beta (Ctrl.) Gamma (Plc.) F-Stat $p_{perm}$ $\eta^{2}$ Result
Mean (STD) Mean (STD) Mean (STD)
Primary Outcome
Naturalness / Human-like Tone 3.955 (0.950) 3.043 (0.825) 3.211 (0.787) 7.017 0.0018 0.187 Sig.
Secondary Interaction Qualities
Memory for Past Talks 3.682 (0.894) 3.348 (0.935) 3.263 (1.147) 1.061 0.3565 0.034 N.S.
Perceived Empathy 3.727 (0.935) 3.435 (1.037) 3.895 (0.737) 1.357 0.2619 0.043 N.S.
Trust (Sharing Feelings) 3.136 (1.390) 2.739 (1.137) 2.737 (1.195) 0.737 0.4907 0.024 N.S.
Personalization 3.136 (1.283) 2.826 (1.114) 3.105 (1.150) 0.463 0.6391 0.015 N.S.
Active Participation 3.636 (1.177) 3.609 (0.891) 3.263 (1.147) 0.751 0.4735 0.024 N.S.
Encouragement to Talk 3.727 (1.202) 3.522 (1.310) 3.105 (1.370) 1.211 0.3025 0.038 N.S.
Off-topic Replies (Lower is better) 1.727 (0.883) 2.304 (1.222) 2.000 (1.247) 1.481 0.2250 0.046 N.S.
Usability and Satisfaction
Overall Satisfaction 3.636 (0.790) 3.348 (0.935) 3.474 (1.124) 0.520 0.6063 0.017 N.S.
Ease of Use 4.000 (1.272) 3.826 (0.984) 3.789 (1.182) 0.204 0.8260 0.007 N.S.
Feature Integration 3.591 (1.008) 3.348 (1.071) 3.737 (0.806) 0.859 0.4463 0.027 N.S.
Predictable / Logical Behavior 3.818 (0.733) 3.826 (0.717) 4.053 (0.911) 0.577 0.5951 0.019 N.S.
Future Use Intention 3.364 (1.177) 2.870 (1.217) 3.316 (1.157) 1.175 0.3243 0.037 N.S.
Appendix D Behavioral and Chat Log Analysis
We performed a quantitative analysis of the chat logs stored in the database to understand the interaction dynamics generated by each architecture. This analysis reveals how the multi-agent FSM architecture (Alpha) fundamentally altered conversational patterns compared to single-agent approaches.
D.1. Interaction Metrics Across Groups
Table 3 summarizes key conversational metrics for all experimental groups. A key finding is the inverse relationship between agent verbosity and perceived naturalness: Alpha produced the highest volume of messages but with the shortest average length, indicating a faster-paced, more conversational rhythm.
Table 3. Conversational Dynamics and Interaction Patterns across Experimental Groups
Metric Alpha (Intervention) Beta (Control) Gamma (Placebo)
Message Volume
Total Agent Messages 459 336 206
Total User Messages 482 336 206
Message Characteristics
Avg. Agent Msg. Length (Chars) 229.8 408.7 575.1
Avg. User Msg. Length (Chars) 29.0 38.9 42.8
Agent:User Length Ratio 7.9:1 10.5:1 13.4:1
D.2. Sentiment Analysis of Chat Content
We performed a lexicon-based sentiment analysis on the message content to evaluate the emotional tone of the agents and users. The analysis used a custom lexicon of positive and negative words in both Persian and English. For each message, we tokenized the text, counted occurrences of positive and negative lexicon terms, and computed a sentiment score as (#pos#neg)/(#pos+#neg)(\#pos-\#neg)/(\#pos+\#neg) when matches were found, or 0 otherwise. Scores range from -1.0 (Negative) to +1.0 (Positive). The Alpha agent exhibited the highest mean sentiment score ($\mu=0.629$), compared to Control ($\mu=0.480$) and Placebo ($\mu=0.465$), aligning with user feedback describing the system as ”warmer” and more supportive. Nearly 80% of Alpha agent messages were classified as positive (Table 4).
Table 4. Sentiment Analysis of Chat Content by Role ($N=66$ participants, all chat messages). Positive/Neutral/Negative classification based on lexicon scoring.
Group Role Mean Score Positive (%) Neutral (%) Negative (%)
Alpha Agent 0.629 79.8 15.0 5.2
User 0.044 13.0 78.1 8.9
Beta Agent 0.480 72.2 18.1 9.7
User 0.055 15.4 74.8 9.7
Gamma Agent 0.465 76.1 11.1 12.8
User 0.010 12.8 75.0 12.2
Appendix E Qualitative Thematic Analysis
Table 5 summarizes the results of the thematic analysis of participants’ open-ended comments, reporting the number of comments in which each theme appeared for each study group. Themes are sentimented and multi-labeled, meaning that a single comment may contribute to multiple themes; positive themes are listed above negative themes for clarity.
Table 5. Theme occurrences by study group (multi-label; a comment may include multiple themes). Positive themes are listed above negative themes.
Theme Description Control Intervention Placebo Total
Positive themes
Engaging Conversation Participants described the interaction as natural, engaging, human-like, or motivating to continue the conversation. 10 9 8 27
Helpful Guidance Participants reported that the chatbot’s suggestions, exercises, or advice were useful, practical, or relevant to their situation. 5 4 4 13
Emotional Support and Accessibility Participants valued ease of access, availability, or the convenience of receiving support without logistical barriers. 2 3 2 7
Empathic Listening Participants emphasized feeling heard, emotionally supported, or comforted by the chatbot’s empathic tone. 1 0 4 5
Negative themes
Interface and Feature Limitations Participants noted missing features or usability issues (like interaction friction, interface limitations) that constrained the experience. 8 6 5 19
Response Delivery and Communication Style Issues Participants criticized response length, fragmentation, clarity, or presentation (such as overly long messages, awkward segmentation). 7 6 5 18
Weak Personalization Participants perceived responses as generic, repetitive, or insufficiently tailored to their context, history, or needs. 5 1 4 10
Conversation Pacing and Question Overload Participants reported poor turn-taking, too many questions at once, topic drift, or pacing that felt rushed or unfocused. 3 2 3 8
Privacy and Safety Concerns Participants raised concerns about data privacy, trust, or the chatbot’s boundaries in sensitive situations. 3 0 3 6